Session: Feb 1999 (session 1)
Wavelength: 0.7 cm


To give feedback on this experiment, click here

Experiment Feedback

Metsahovi (Kirsi Karlamaa at Thu Mar 4 07:49:36 MET 1999) Minor failures
Clouds. Problems with BBCs, IF-level might have been too low.

Effelsberg (dgraham at Thu Mar 4 09:23:10 MET 1999) Minor failures
Pointing probably poor. Due to reduced azimuth speed (18deg/min) lost most calibrator scans. Lost 0200-0240 with tape and antenna problems.

OAN\-Yebes (Paco Colomer at Thu Mar 4 10:58:42 MET 1999) Success
All scans observed, without known technical problems. Weather poor during first half of experiment. Non standard Tsys measurement - more info in log.

Medicina (A. Cattani at Thu Mar 4 11:20:19 MET 1999) Minor failures
No problems regarding acquistion rack and receiver. Problems with pointing that shows poor.

Calibration (Jean\-Francois Desmurs at Mon Mar 8 11:44:46 MET 1999) Success
The ANTAB formatted file for this experiment can be found on VLBEER in the feb99 directory under the name: gj009.antab. Read the comments in the file to get further information.

Onsala (A. Polatidis at Mon Mar 8 13:10:21 MET 1999) Minor failures
From start until 2011 UT, the VC were set at wrong frequencies because a wrong LO value was used in the schedule. No other problems after the VC frequiencies were fixed. Rain during most of the experiment.